OpenStudy (anonymous):

Please help :( The slope of line JD is 1/4, J has coordinates (10, -4), and D has coordinates (6, y). What is the value for y?

OpenStudy (anonymous):

m = Δy/ Δx ( y + 4) / ( 6 - 10 ) = 1/4 => y = ...? Can you fill in the blank

OpenStudy (anonymous):

well, i got as far as to: y + 4 / -4 = 1/4 but I'm not positive what to do after that.

OpenStudy (anonymous):

-> y + 4 = -1 Can you finish it now?

OpenStudy (anonymous):

how'd you get that?

OpenStudy (anonymous):

-4/4 = -1

OpenStudy (anonymous):

and how'd you get that?

OpenStudy (anonymous):

like what exactly do I have to do from: y-4/-4 = 1/4?

OpenStudy (anonymous):

oh, so y would be -5?

OpenStudy (anonymous):

wouldn't it be y +4 = -1?

OpenStudy (anonymous):

Yep, x = -5 !

OpenStudy (anonymous):

Thank you SO much x) you helped me a lot :)

OpenStudy (anonymous):

Thanks xD I'm a very open person so I just ask what I need to ask LOL I hate not knowing or understanding things x)

OpenStudy (anonymous):

LOL I know right? It's so frustrating sometimes when people do so xD and thanks x)

OpenStudy (anonymous):

Sometime it's tough to report to god that I've done a good deed today :P

OpenStudy (anonymous):

LOL I'll send an extra thought to him for you ;) LOL

OpenStudy (anonymous):

Nice =)

OpenStudy (anonymous):

hey, can you help me with another geometry question? x)